# PULSE: Practical Evaluation Scenarios for Large Multimodal Model Unlearning

### v1

@@ 2 Related Work
Target Model Unlearning of Fine-Tuned Knowledge Unlearning of Pre-trained Knowledge Sustainability
@@ 3 PULSE Protocol
In addition to the typical unlearning evaluation pipeline that first fine-tunes the unlearning target samples and then conduct unlearning on it, our PULSE evaluates (i) unlearning on pre-trained knowledge and (ii) sustainability against multiple unlearning requests. In this section, we introduce the general problem setup (Section 3.1) and the detailed setup of evaluation on each perspective (Sections 3.2, 3.3 and 3.4).
@@ 3.4 Sustainability
The importance of such evaluation axis is first proposed by MUSE [9] for LLMs; here, we extend it to the multimodal setting for LMMs. Figure 1(b) illustrates the pipeline of sustainability evaluation. In contrast to the single-step unlearning in Section 3.2, here we divide $\mathcal{D}_{\text{unlearn}}$ into several subsets and make the target model unlearn them sequentially. During this process, we track how the model’s generality and effectiveness change after each operation.
@@
Table 2: Comparison of Experimental Settings. The data used in the experiments were selected from the publicly released MLLMU-Bench [10] dataset to match our experimental configurations. Each individual is associated with 10 questions—half text-only and half multimodal. Therefore, the dataset size equals the number of individuals multiplied by 10.
Type of Knowledge to Unlearn Number of Individuals in $\mathcal{D}_{\text{unlearn}}$ Unlearning Count Individuals Unlearned per Operation

### v2

@@ 2 Related Work
Target Model
Unlearning of
Fine-Tuned Knowledge
Unlearning of
Pre-trained Knowledge
Sustainability
@@ 3 PULSE Protocol
In addition to the typical unlearning evaluation pipeline that first fine-tunes the unlearning target samples and then conduct unlearning on it, our PULSE evaluates (i) unlearning on pre-trained knowledge and (ii) sustainability against multiple unlearning requests. In this section, we introduce the general problem setup (Section˜3.1) and the detailed setup of evaluation on each perspective (Sections˜3.2, 3.3 and 3.4).
@@ 3.4 Sustainability
The importance of such evaluation axis is first proposed by MUSE [9] for LLMs; here, we extend it to the multimodal setting for LMMs. Figure 1(b) illustrates the pipeline of sustainability evaluation. In contrast to the single-step unlearning in Section˜3.2, here we divide $\mathcal{D}_{\text{unlearn}}$ into several subsets and make the target model unlearn them sequentially. During this process, we track how the model’s generality and effectiveness change after each operation.
@@
[…]
Type of Knowledge
to Unlearn
Number of Individuals
in $\mathcal{D}_{\text{unlearn}}$
Unlearning Count
Individuals Unlearned
per Operation